Solve the following equations:
 (5/2) x = -5..........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

First we have to multiply both the side by 2,
= (5x/2) × 2 = - 5 × 2
= 5x = - 10
Now,
We have to divide both the side by 5,
Then we get,
= 5x/5 = -10/5
= x = -2

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
(5/2)x = -5
5x / 2 = -5
5x = -10
x = -10/5
x = -2
 
Hope it helps!
Thank you ✌️